The number of mosquitoes m(x) in millions, in a certain area depends on the june rainfall x, in inches. m(x)-20x-x^2. What rainfall produces the max number of mosquitoes? 400, 10,0,20
Answer by mukhopadhyay(490)

m(x)= 20x-x^2
m(x) represents a parabola, opens downward (because of negative sign in x^2)
The maximum will be attained at its vertex.
x-coordinate of the vertex of a parabola ax^2 + bx + c is -b/2a
-b/2a = -20/(-2 ) = 10
Answer: 10